### v1

@@ 4.3 NLTE effects on Eu II lines
Fig. 4 shows the NLTE corrections for Eu II 4129Å line (left panel) and 6645Å lines (right panel) plotted against the metallicity [Fe/H] of the model atmospheres. The red line depicts the NLTE correction for red giants (RG) with parameters of =4500/=1.5, while the black line illustrates the NLTE correction for main-sequence stars (MS) with parameters of =6000/=4.0. Additionally, the blue line represents the NLTE correction of subgiant (SG) with parameters of 5500/3.5. The NLTE corrections for Eu II 4129 Å are positive, while negative or close to zero for Eu II 6645 Å . For Eu II 4129 Å line, at values of [Fe/H] near dex, the corrections for RG, SG, and MS are close to each other. For Eu II 6645 Å line, the NLTE corrections for RG are slightly higher than for SG, and the corrections for SG are slightly higher than those for MS stars when [Fe/H] is less than 0 dex. For the MS, the NLTE corrections do not exceed dex at [Fe/H] = dex. However, for RG, the NLTE corrections can reach dex at [Fe/H]= dex. It also indicates that the NLTE corrections increase with decreasing metallicity, which may have an impact on the GCE trend of [Eu/Fe] especially with the lower [Fe/H] (Alexeeva et al. 2023; Storm et al. 2024). Although the NLTE corrections for the Eu II 4129 Å and Eu II 6645 Å lines are entirely opposite, the final results after NLTE corrections match (as discussed in Section 4.4), further proving the necessity of NLTE corrections for determination of Eu abundance.
@@
We show the departure coefficients, , for Eu II 6645 Å line as a function of optical depth in Fig. 5. Here, the departure coefficient = is defined as the ratio of the NLTE population of atomic level to the LTE population of atomic level . The solid lines represent the lower levels, while the dashed lines represent the upper levels. The departure coefficients for the solar model atmosphere are depicted in red, while the black one represents the model atmosphere with K, dex, and [Fe/H] = dex. We chose this model atmosphere because its parameters are similar to those of most of our sample stars. The formation height is shown as + at log() = 0 for line center. In the solar atmosphere, the formation height of 6645 Å line corresponds to log() = 0.64, where both states and tend to be overpopulated compared to LTE values, with being higher than at the same time. According to Bergemann & Nordlander (2014a), this suggests that the line source function exceeds the Planck function (/ / ). As a result, line became weaker under NLTE conditions relative to LTE. This ultimately causes slightly positive NLTE corrections. For the model atmosphere with K, dex, and [Fe/H] = dex, another effect plays a more important role. The departure coefficient of upper energy state is closer to that of lower energy state () at log() = 0.6. Thus, the enhanced line opacity compared to LTE (since 1), increases the number of absorbed photons. This enhanced absorption strengthens the line under NLTE conditions, leading to a negative NLTE correction.
@@
We present the fitting results of four stars as examples in Fig. 6. For Eu II 4129 Å line, We present the observed spectrum as black dots in the upper left panel with parameters of =4852 K, =2.12 dex, [Fe/H]= dex and in the lower left panel with =4988 K, =2.53 dex, [Fe/H]= dex. For Eu II 6645 Å line, we show the observed spectrum in black dots with parameters of =4995 K, =2.42 dex, [Fe/H]= dex in the upper right panel and =4401 K, =1.13 dex, [Fe/H]= dex in the lower right panel. The red and blue lines represent the best-fit synthetic spectra based on 1D LTE and NLTE models, respectively. We show the comparison of Eu/Fe differences between the 4129 Å and 6645 Å lines in Fig. 7. The NLTE corrections result in more consistent values between the two lines, highlighting the importance of NLTE corrections for precise Eu abundance determinations.
@@ 4.4 Comparison of the results with GCE model
This table is available in its entirety in machine-readable form. The first five entries are shown here for guidance regarding its format and content.
@@
This table is available in its entirety in machine-readable form. The first five entries are shown here for guidance regarding its format and content.
@@
The wavelength range for some of our stars does not include the Eu II 4129 Å line, and in some cases, both the Eu II 4129 Å and Eu II 6645 Å lines are blended or very weak. Therefore, the results presented below are only for the spectra that were well fitted with a total of 164 stars having metallicities ranging from to dex. However, the 3D grid covers a smaller range of atmospheric parameters (Magic et al. 2013a, b) and thus only a small part of our sample of stars was fitted with these models. Therefore, our results are only based on the results of the 1D. All the best fitting results are provide in Tab. 3 for Eu II 4129 Å and Tab. 4 for Eu II 6645 Å . There are five stars in our sample that are carbon-enhanced metal-poor with s-process element enhancement (CEMP-s) stars (Placco et al. 2018), which do not follow the Galactic chemical evolution trend, and we use * to mask them in Table 3 and 4 .
@@
Fig. 8 shows the [Eu/Fe] based on 1D NLTE of our metal-poor sample as a function of [Fe/H] combined with 1D LTE Eu abundance measurements of 1274 metal-rich stars from Gaia-ESO survey (Gilmore et al. 2022; Randich et al. 2022) with high spectral signal-to-noise ratio (SNR; Lian et al. 2023). We divided the bins to ensure approximately an equal number of stars in each, with the green squares and blue dots representing the averaged NLTE and LTE [Eu/Fe] ratios across the selected bins. The yellow triangles and cyan circles represent the NLTE and LTE [Eu/Fe] values for a total of 141 stars based on the 4129 Å line and 35 stars based on the 6645 Å line combined. As shown in Fig. 8, the trend of [Eu/Fe] is almost flat in metal-poor region and both the scatter of data points and the NLTE adjustments tend to decrease from low to high metallicity. However, this phenomenon requires confirmation with more data points in the future.
@@
For comparison, the tracks of two galactic chemical evolution (GCE) models generated by OMEGA+ Côté et al. (2018b) using the parameters of the basic model in Lian et al. (2023) are presented. In these models, to account for the rapid decrease of [Eu/Fe] in metal-rich regime which implies much shorter release timescale of Eu than Fe by SN-Ia, the contribution of Eu in short timescale from magneto rotating supernova (MRSN) has been taken into account in addition to the neutron star merger (NSM) which releases Eu in much longer timescales (Côté et al. 2019). For NSM we assume an ejecta mass of 2.5 with yields table from Arnould et al. (2007) and occurance rate of 2 for every solar mass of star formation. The delay time distribution of NSM is assumed to follow a power law in the form of with minimum delay time of 10 Myr and maximum delay time of 106 Gyr. To include the contribution of MRSN, we replace a tiny fraction (0.013% and 0.015%) of core-collapse supernova (CCSN) in the mass range of 13-25 by MRSN and use the yields table from Nishimura et al. (2015). This fraction is higher than that (0.01%) adopted in Lian et al. (2023) in which the data in the low metallicity regime is limited and the fraction of MRSN is not well constrained. In this work we intend to test the impact of the NLTE corrections on the required parameters responsible for Eu production in GCE models. For simplicity, we focus on one of the key and poor constrained parameter, the fraction of MRSN in CCSN (), which is positively correlated with the [Eu/Fe] abundance in old metal-poor stars. As expected, with higher [Eu/Fe] after NLTE correction, a GCE model with larger is needed to well fit the average [Eu/Fe] of our sample. We note that the required change in () is not substantial given the NLTE correction derived in this work.
@@ 5 Summary
Eu, as an r-process element, provides insight into violent events such as neutron star mergers, helping us to better understand nucleosynthetic production sites. Due to the lack of observational data for r-process elements, especially for Eu based on NLTE analysis in the metal-poor region, this work aims to improve the understanding of Eu abundances in metal-poor stars by applying NLTE and 3D corrections.
We performed a detailed analysis of the NLTE effects on Eu II 4129 Å and Eu II 6645 Å for a sample of metal-poor stars. For Eu II 4129 Å line, we found that the NLTE effects result in positive NLTE abundance correction in both 1D and 3D in solar atmosphere. For Eu II 6645 Å line, we found that the NLTE effects result in a slightly positive NLTE abundance correction in 1D and a negative one in 3D in solar atmosphere. Although the NLTE corrections for the Eu II 4129 Å and Eu II 6645 Å lines are entirely opposite, the discrepancy between the abundances derived from individual lines decreases after NLTE fitting, once again showcasing the importance of NLTE abundance determination. Finally, we show the trend of [Eu/Fe] as a function of [Fe/H] and compare it with the GCE models. It indicates that the NLTE correction does not require significant change of the parameter for the Eu production in GCE models. However, due to the limitations of observational data and models, this phenomenon requires confirmation with more data points, e.g. 4MOST in the future.
@@ Acknowledgements.
This work is supported by the Natural Science Foundation of China (Nos. 12288102,12125303,12090040/3,12103064,12403039,12373036), the National Key R&D Program of China (grant Nos. 2021YFA1600403/1, 2021YFA1600400), and the Natural science Foundation of Yunnan Province (Nos. 202201BC070003, 202001AW070007), the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001) and the “Yunnan Revitalization Talent Support Program”-Science and Technology Champion Project (N0. 202305AB350003). NS and MB acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949173). MB is supported through the Lise Meitner grant from the Max Planck Society. H.-L.Y. acknowledges support from the Youth Innovation Promotion Association of the CAS. We acknowledge support by the Collaborative Research centre SFB 881 (projects A5, A10), Heidelberg University, of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).References

### v2

@@ 4.3 NLTE effects on Eu II lines
Figure 4 shows the NLTE corrections for the Eu II 4129Å line (left panel) and 6645Å lines (right panel) plotted against the metallicity [Fe/H] of the model atmospheres. The red line depicts the NLTE correction for red giants (RG) with parameters of =4500/=1.5, while the black line illustrates the NLTE correction for main-sequence stars (MS), with parameters of =6000/=4.0. Additionally, the blue line represents the NLTE correction of subgiant (SG) with parameters of 5500/3.5. The NLTE corrections for Eu II 4129 Å are positive, while negative or close to zero for Eu II 6645 Å . For Eu II 4129 Å line, at values of [Fe/H] near dex, the corrections for RG, SG, and MS are close to each other. For the Eu II 6645 Å line, the NLTE corrections for RG are slightly higher than for SG, and the corrections for SG are slightly higher than those for MS stars when [Fe/H] is less than 0 dex. For the MS, the NLTE corrections do not exceed dex at [Fe/H] = dex. However, for RG, the NLTE corrections can reach dex at [Fe/H]= dex. It also indicates that the NLTE corrections increase with decreasing metallicity, which may have an impact on the GCE trend of [Eu/Fe] especially with the lower [Fe/H] (Alexeeva et al. 2023; Storm et al. 2024). Although the NLTE corrections for the Eu II 4129 Å and Eu II 6645 Å lines are entirely opposite, the final results after NLTE corrections match (as discussed in Section 4.4), further proving the necessity of NLTE corrections for determination of Eu abundance.
@@
We show the departure coefficients, , for Eu II 6645 Å line as a function of optical depth in Fig. 5. Here, the departure coefficient, =, is defined as the ratio of the NLTE population of atomic level to the LTE population of atomic level, . The solid lines represent the lower levels, while the dashed lines represent the upper levels. The departure coefficients for the solar model atmosphere are depicted in red, while the black one represents the model atmosphere with K, dex, and [Fe/H] = dex. We chose this model atmosphere because its parameters are similar to those of most of our sample stars. The formation height is shown as + at log() = 0 for line center. In the solar atmosphere, the formation height of 6645 Å line corresponds to log() = 0.64, where both states and tend to be overpopulated compared to LTE values, with being higher than at the same time. According to Bergemann & Nordlander (2014a), this suggests that the line source function exceeds the Planck function (/ / ). As a result, the line becomes weaker under NLTE conditions relative to LTE. This ultimately causes slightly positive NLTE corrections. For the model atmosphere with K, dex, and [Fe/H] = dex, another effect plays a more important role. The departure coefficient of upper energy state is closer to that of lower energy state () at log() = 0.6. Thus, the enhanced line opacity compared to LTE (since 1), increases the number of absorbed photons. This enhanced absorption strengthens the line under NLTE conditions, leading to a negative NLTE correction.
@@
We present the fitting results of four stars as examples in Fig. 6. For the Eu II 4129 Å line, the observed spectrum is shown as black dots in the upper left panel, with parameters: =4852 K, =2.12 dex, and [Fe/H]= dex; in the lower left panel, we have: =4988 K, =2.53 dex, [Fe/H]= dex. For the Eu II 6645 Å line, we show the observed spectrum in black dots with parameters of =4995 K, =2.42 dex, and [Fe/H]= dex in the upper right panel and =4401 K, =1.13 dex, and [Fe/H]= dex in the lower right panel. The red and blue lines represent the best-fit synthetic spectra based on 1D LTE and NLTE models, respectively. We show the comparison of [Eu/Fe] differences between the 4129 Å and 6645 Å lines in Fig. 7. The NLTE corrections result in more consistent values between the two lines, highlighting the importance of NLTE corrections for precise Eu abundance determinations.
@@ 4.4 Comparison of the results with GCE model
This table is available in its entirety in machine-readable form at the CDS.
@@
This table is available in its entirety in machine-readable form at the CDS.
@@
The wavelength range for some of our stars does not include the Eu II 4129 Å line, and in some cases, both the Eu II 4129 Å and Eu II 6645 Å lines are blended or very weak. Therefore, the results presented below are only for the spectra that were well fitted with a total of 164 stars having metallicities ranging from to dex. However, the 3D grid covers a smaller range of atmospheric parameters (Magic et al. 2013a, b) and, thus, only a small proportion of our sample of stars was fitted with these models. Therefore, our results are only based on the results of the 1D. All the best fitting results are provide in Table 3 for Eu II 4129 Å and Table 4 for Eu II 6645 Å . There are five stars in our sample that are carbon-enhanced metal-poor with s-process element enhancement (CEMP-s) stars (Placco et al. 2018), which do not follow the Galactic chemical evolution trend. We use * to mask them in Tables 3 and 4 .
@@
In Fig. 8, we show the [Eu/Fe] based on 1D NLTE of our metal-poor sample as a function of [Fe/H] combined with 1D LTE Eu abundance measurements of 1274 metal-rich stars from Gaia-ESO survey (Gilmore et al. 2022; Randich et al. 2022) with a high spectral signal-to-noise ratio (S/N; Lian et al. 2023). We divided the bins to ensure approximately an equal number of stars in each, with the green squares and blue dots representing the averaged NLTE and LTE [Eu/Fe] ratios across the selected bins. The yellow triangles and cyan circles represent the NLTE and LTE [Eu/Fe] values for a total of 141 stars based on the 4129 Å line and 35 stars based on the 6645 Å line combined. As shown in Fig. 8, the trend of [Eu/Fe] is almost flat in metal-poor region and both the scatter of data points and the NLTE adjustments tend to decrease from low to high metallicity. However, this phenomenon requires confirmation based on more data points in the future.
@@
For comparison, the tracks of two galactic chemical evolution (GCE) models generated by OMEGA+ Côté et al. (2018b) using the parameters of the basic model in Lian et al. (2023) are presented. In these models, to account for the rapid decrease of [Eu/Fe] in the metal-rich regime which implies much shorter release timescale of Eu than Fe by SN-Ia, the contribution of Eu on short timescale from magneto-rotating supernova (MRSN) has been taken into account, in addition to the neutron star merger (NSM) that releases Eu on much longer timescales (Côté et al. 2019). For NSM, we assumed an ejecta mass of 2.5 with yields table from Arnould et al. (2007) and occurrence rate of 2 for every solar mass of star formation. The delay time distribution of NSM is assumed to follow a power law in the form of with minimum delay time of 10 Myr and maximum delay time of 106 Gyr. To include the contribution of MRSN, we replaced a tiny fraction (0.013% and 0.015%) of core-collapse supernova (CCSN) in the mass range of 13-25 by MRSN and used the yields table from Nishimura et al. (2015). This fraction is higher than that (0.01%) adopted in Lian et al. (2023), where the data in the low-metallicity regime is limited and the fraction of MRSN is not well constrained. In this work we intend to test the impact of the NLTE corrections on the required parameters responsible for Eu production in GCE models. For simplicity, we focus on one of the key and poor constrained parameter, the fraction of MRSN in CCSN (), which is positively correlated with the [Eu/Fe] abundance in old metal-poor stars. As expected, with higher [Eu/Fe] after NLTE correction, a GCE model with larger is needed to well fit the average [Eu/Fe] of our sample. We note that the required change in () is not substantial given the NLTE correction derived in this work.
@@ 5 Summary
As an r-process element, Eu provides insights into violent events such as neutron star mergers, helping us to improve our understanding of nucleosynthetic production sites. Due to the lack of observational data for r-process elements, especially for Eu, based on NLTE analysis in the metal-poor region, this work is aimed at investigating Eu abundances in metal-poor stars by applying NLTE and 3D corrections.
We performed a detailed analysis of the NLTE effects on Eu II 4129 Å and Eu II 6645 Å for a sample of metal-poor stars. For the Eu II 4129 Å line, we found that the NLTE effects result in positive NLTE abundance correction in both 1D and 3D in the solar atmosphere. For the Eu II 6645 Å line, we found that the NLTE effects result in a slightly positive NLTE abundance correction in 1D and a negative one in 3D in the solar atmosphere. Although the NLTE corrections for the Eu II 4129 Å and Eu II 6645 Å lines are entirely opposite, the discrepancy between the abundances derived from individual lines decreases after NLTE fitting, once again showcasing the importance of NLTE abundance determination. Finally, we show the trend of [Eu/Fe] as a function of [Fe/H] and compare it with the GCE models. It indicates that the NLTE correction does not require significant change of the parameter for the Eu production in GCE models. However, due to the limitations of observational data and models, confirming this phenomenon would require more data points, for instance, with 4MOST in the future.
@@ Acknowledgements.
This work is supported by the Natural Science Foundation of China (Nos. 12288102,12125303,12090040/3,12103064,12403039), the National Key R&D Program of China (grant Nos. 2021YFA1600403/1, 2021YFA1600400), and the Natural science Foundation of Yunnan Province (Nos. 202201BC070003, 202001AW070007), the International Centre of Supernovae, Yunnan Key Laboratory (No. 202302AN360001) and the “Yunnan Revitalization Talent Support Program”-Science and Technology Champion Project (N0. 202305AB350003). NS and MB acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 949173). MB is supported through the Lise Meitner grant from the Max Planck Society. We sincerely thank Dr. Hongliang Yan for his valuable suggestions that improved the quality of this paper. We acknowledge support by the Collaborative Research centre SFB 881 (projects A5, A10), Heidelberg University, of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).References